In the morning Maria drove to a business appointment at 40 mph. Her average speed on the return trip in the afternoon was 30 mph. The trip took 1/3 hour longer because of heavy traffic. How far did she travel to the appointment?
Answer by abdulrasid84@gmail.com(4)

we know that Time is equal Distance over Rate T=D/R
Let D is the Distance.
And Let we say. time = t1 + t2
time = t1 + t2
1/3 = d/40 + d/30
then multiply both side by 40.
13.33 = 40d/40 + 40d/30
13.33 = d + 1.33d
13.33 = 2.33d
then divide 13.33 by 2.33
d = 5.721
Therefor maria did travel at 5.721 m.